# Artificial Intelligence AUC Area Under the Curve

### v2

@@ Appendix A Limitations and Societal Impact
This work opens the door to exploring deterministic initializations that exploit symmetry to endow neural networks with desirable properties beyond the reach of stochastic schemes. Among the possible designs, we focused on the Sinusoidal initialization due to its natural structure and analytical tractability. Nevertheless, the broader space of structured deterministic initializations remains largely unexplored, offering a fertile avenue for future research.
@@
Regarding the societal impact, the proposed Sinusoidal initialization enables faster convergence during training, reducing the number of epochs required to reach target performance. This directly translates into lower computational demands, which in turn reduces both the economic cost and the environmental footprint of training deep neural networks.
@@ Appendix B Theoretical details
This appendix provides the theoretical foundations underlying the results stated in Section 3, and in particular formal proofs of Theorems 1 and 2. We also present a general version of Theorem 1, extending the simplified setting used in the main text to a more flexible framework that covers non-identically distributed inputs and milder moment assumptions.
@@ B.2 Row cancellation in the Sinusoidal initialization
Let $W\in\mathbb{R}^{m\times n}$ be defined as in Equation (1). Then, for every row index $i\in\{1,\dots,m\}$, the entries along that row sum to zero:
@@
To show this, let $\omega\in\mathbb{C}$ be the unitary complex number and angle $2\pi i/n$. It is clear that $w^{n}=1$, and hence
@@ Appendix C Experimental Setup
For full reproducibility, all code used to conduct the experiments is publicly available in the accompanying GitHub repository: https://github.com/.
@@ C.3 Benchmark Evaluation
This subsection provides the full experimental details for the training process corresponding to the experiments presented in Section 4, and summarized in Table 3 and Figures 1, 6, and 9. The tables below detail the training hyperparameters (Table 4) and hardware (Table 5) used for each model-dataset pair.
@@
ResNet-50 CIFAR-100 100 SGD, Adam, AdamW $10^{-3}$ $10^{-3}$ 64
@@ D.1 Arcsine random initialization
We replicated the experiments from Figure 4 and Figure 5 using this randomized initialization. The results, shown in Figure 8, demonstrate that this initialization behaves similarly to other stochastic methods: the activations exhibit comparable neuron skewness. This confirms that the stochastic nature of the weights, not the distribution, is responsible for the observed effects.
These findings make it clear that the structure introduced by the Sinusoidal initialization is the critical component driving its benefits. This outcome is consistent with the theory established in Subsection 3.2, and further reinforces that the deterministic layout of weights plays a central role in mitigating activation skewness.
@@
D.2 Additional Training Plots

### v3

@@ Appendix A Limitations and Societal Impact
This work opens the door to exploring deterministic initializations that exploit symmetry to endow neural networks with desirable properties beyond the reach of stochastic schemes. Among the possible designs, we focused on the Sinusoidal initialization due to its natural structure and analytical tractability. Nevertheless, the broader space of structured deterministic initializations remains scarcely explored, offering a fertile avenue for future research.
@@
Regarding the societal impact, the proposed Sinusoidal initialization enables faster convergence during training, reducing the number of epochs required to reach target performance. This directly translates into lower computational demands, which in turn reduces both the economic cost and the environmental footprint of training DNNs.
@@ Appendix B Theoretical details
This appendix provides the theoretical foundations underlying the results stated in Section 3, and in particular formal proofs of Theorems 1, 2 and  3. We also present a general version of Theorem 1, extending the simplified setting used in the main text to a more flexible framework that covers non-identically distributed inputs and milder moment assumptions.
@@ B.1 Threshold equivalence
Theorem ([Theorem 2 (Asymptotic skewness under random initialization)).
] Consider a neuron with zero bias and input $X_{1},\ldots,X_{n}$ and weights $W_{1},\ldots,W_{n}$. Assume $\{W_{i}\}$ are i.i.d. with $\mathbb{E}[W_{1}]=0$ and $\operatorname{Var}(W_{1})=\theta^{2}\in(0,\infty)$, and $\{X_{i}\}$ are i.i.d., independent of $\{W_{i}\}$, with $\mathbb{E}[X_{1}]=\mu>0$ and $\operatorname{Var}(X_{1})=\sigma^{2}<\infty$. Let $Z_{n}=\sum_{i=1}^{n}W_{i}X_{i}$. Then for every deterministic sequence $\alpha_{n}\downarrow 0$, the probability of having
$\bigl|P(Z_{n}>0\mid W_{1},\ldots,W_{n})-1/2\bigr|>\alpha_{n}$
tends to one as $n\to\infty$. Equivalently, with probability tending to one there exists a nonzero skewness level at which the neuron is skewed.
Proof.
Write $S_{n}=\sum_{i=1}^{n}W_{i}$ and fix any sequence $\alpha_{n}\downarrow 0$. By the Threshold Equivalence Theorem 1, for each fixed $\alpha\in(0,1/2)$ the event
$E_{n}(\alpha):=\Bigl\{\bigl|\mathbb{P}(Z_{n}>0\mid W_{1},\ldots,W_{n})-1/2\bigr|>\alpha\Bigr\}$
is, with probability tending to one, equivalent to
$F_{n}(\alpha):=\Bigl\{|S_{n}|>\lambda_{n}(\alpha)\Bigr\},$
where
$\lambda_{n}(\alpha)=(\theta\sigma\sqrt{n})/\mu\cdot\Phi^{-1}\!\bigl(1/2+\alpha\bigr).$
To apply this along a vanishing sequence, choose a decreasing sequence $\alpha^{(k)}\downarrow 0$ and integers $N_{k}\uparrow\infty$ such that, for all $n\geq N_{k}$, the difference between $E_{n}(\alpha^{(k)})$ and $F_{n}(\alpha^{(k)})$ has probability at most $2^{-k}$. Define a piecewise constant selection $\alpha_{n}=\alpha^{(k)}$ for $N_{k}\leq n<N_{k+1}$. This ensures
$\mathbb{P}\bigl(E_{n}(\alpha_{n})\triangle F_{n}(\alpha_{n})\bigr)\longrightarrow 0.$
It remains to show that $\mathbb{P}\bigl(F_{n}(\alpha_{n})\bigr)\to 1$. By the Central Limit Theorem,
$\frac{S_{n}}{\theta\sqrt{n}}\Rightarrow\mathcal{N}(0,1).$
Since $\alpha_{n}\to 0$ and $\Phi^{-1}$ is continuous at $1/2$, we have that
$c_{n}:=\frac{\lambda_{n}(\alpha_{n})}{\theta\sqrt{n}}=\frac{\sigma}{\mu}\,\Phi^{-1}\!\bigl(\tfrac{1}{2}+\alpha_{n}\bigr)\longrightarrow 0.$
Thus, if $\mathcal{N}(0,1)$ is denoted by $G$, it follows that
$\mathbb{P}\bigl(F_{n}(\alpha_{n})\bigr)=\mathbb{P}\!\left(\left|\frac{S_{n}}{\theta\sqrt{n}}\right|>c_{n}\right)\longrightarrow\mathbb{P}(|G|>0)=1.$
Combining both steps yields
$\mathbb{P}\bigl(E_{n}(\alpha_{n})\bigr)\geq\mathbb{P}\bigl(F_{n}(\alpha_{n})\bigr)-\mathbb{P}\bigl(E_{n}(\alpha_{n})\triangle F_{n}(\alpha_{n})\bigr)\longrightarrow 1,$
and the result follows. ∎
@@ B.2 Row cancellation in the Sinusoidal initialization
Let $W\in\mathbb{R}^{m\times n}$, with $m<n$, be defined as in Equation (1). Then, for every row index $i\in\{1,\dots,m\}$, the entries along that row sum to zero:
@@
To show this, let $\omega\in\mathbb{C}$ be the unitary complex number and angle $2\pi i/n$, which clearly satisfies that $w^{n}=1$. As $i\leq m<n$, it follows that $\omega\neq 1$, and hence
@@ Appendix C Experimental Setup
For full reproducibility, all code used to conduct the experiments is publicly available in the accompanying GitHub repository: https://github.com/jmiravet/Sinusoidal-Initialization.
@@ C.3 Benchmark Evaluation
This subsection provides the full experimental details for the training process corresponding to the experiments presented in Section 4, and summarized in Table 3 and Figures 1, 6, and 11. The tables below detail the training hyperparameters (Table 4) and hardware (Table 5) used for each model-dataset pair.
@@
ResNet‑50 CIFAR‑100 100 SGD, Adam, AdamW $10^{-3}$ $10^{-3}$ 64
@@ D.1 Arcsine random initialization
We replicated the experiments from Figure 4 and Figure 5 using this randomized initialization. The results, shown in Figure 8, demonstrate that this initialization behaves similarly to other stochastic methods: the activations exhibit comparable neuron skewness. This confirms that the stochastic nature of the weights, not the distribution itself, is responsible for the observed effects.
@@
These findings make it clear that the structure introduced by the Sinusoidal initialization is the critical component driving its benefits. This outcome is consistent with the theory established in Subsection 3.2, and further reinforces that the deterministic layout of weights plays a central role in mitigating activation skewness. To further support this conclusion, we replicated the experiment reported in Section 4 on ResNet-50 with CIFAR-100 and SGD. The training results, presented in Figure 9, clearly show how the randomness of the arcsine variant delays convergence with respect to the structured Sinusoidal initialization, which provides a faster training process.
Figure 9: Training curves of the Sinusoidal initialization and its randomized counterpart, showing that the structural arrangement of weights accelerates the initial stages of training.
D.2 Comparison with other deterministic initializations
In addition to stochastic baselines, a few deterministic initialization schemes have also been proposed in the literature, namely those introduced in zhao2022zero and blumenfeld2020beyond, already referenced in Section 2. To illustrate the relative performance of these methods, we provide a representative experiment on ResNet-50 with CIFAR-100 using SGD, under the same hyperparameter configuration described in Section 4. While this comparison is restricted to a single training setup, it provides a direct assessment of how these deterministic strategies behave under identical conditions. The results are shown in Figure 10.
Figure 10: Training curves of Default, Sinusoidal, Zero and Blumenfeld initializations, showing that our approach consistently outperforms the other deterministic methods.
As can be observed, the Sinusoidal initialization clearly outperforms both alternatives, achieving faster convergence and higher final accuracy. This experiment thus provides compelling evidence that, within the family of deterministic approaches, our method offers a distinct advantage.
D.3 Additional Training Plots